An urn contains 1 green ball, 1 red ball, 1 yellow ball and 1 white ball. I draw 4 balls with replacement. What is the probability that there is at least one color that is appears exactly twice?
Answer #1

probability that there is at least one color that is appears exactly twice

=P(select two colors from 4 and both appear twice and arrange them)+P(select a color from 4 which appear twice, select 2 out of 4 places for it and then arrange two different colors in 2 remaining places)

=(4C2)*(4C2)/44 +(4C1)*(4C2)*3*2/44 =6*6/256+4*6*3*2/256 =180/256 =45/64